# Cryptomite: A versatile and user-friendly library of randomness extractors

### v1

@@ 3.2.1 Two-source extractors in the product sources model
Any strong classical-proof (or quantum-proof) $(k_{1},n_{2},\epsilon,n_{1},n_{2},m)$-seeded extractor is a strong classical-proof (or quantum-proof) $(k_{1},k_{2},2^{n_{2}-k_{2}}\epsilon,n_{1},n_{2},m)$-two-source extractor, strong in the (now) weak seed. If the extractor is constructed from a two-universal family of hash functions, it is a strong classical-proof (or quantum-proof) $(k_{1},k_{2},2^{(n_{2}-k_{2})/2}\epsilon,n_{1},n_{2},m)$-two-source extractor.
The quantum-proof claim in Theorem 1 does not directly cover quantum side-information about the (now) weak seed, only on the weak input, and still demands that the weak seed and weak input are independent. Using the proof techniques of [36] (‘Proof of proposition 1’) this can be directly generalised to the case that an adversary has quantum side-information about both the input and the weak seed, but with the constraint that, for extractors based on two-universal hash families, this side-information is in product form. More concretely, this is the case when the input and the now weak seed form a so-called classical-classical quantum-quantum state, whereby the combined two-universal hashing based extractor input (for the input and (weak) seed given by Equation 4) takes the form
@@
This model for quantum-proof extractors is the product sources model.
Theorem 2 (Quantum-proof two-source extension in the product sources model, Theorem 1 with Proposition 1 in [36]).
Any strong quantum-proof $(k_{1},n_{2},\epsilon,n_{1},n_{2},m)$-seeded extractor based on two-universal hash families is a strong quantum-proof $(k_{1},k_{2},2^{(n_{2}-k_{2})/2}\epsilon,n_{1},n_{2},m)$-two-source extractor in the product sources model.
@@
The proof of this follows directly from [36] ‘Proof of proposition 1’ by noticing that $n-1$ can be generically replaced with $n_{2}$, the length of the (weak) seed and that their bound on (*)(*)( * ) (as appearing in the text) is found by bounding the collision entropy of the seeded version of the extractor, which holds for all extractors based off two-universal families of hash functions. ∎
@@ 3.2.2 Two-source extractors in the Markov model
Two-source extractors can also be made secure in the so-called Markov model [11], which considers evens stronger adversaries than in the product sources model. Instead of requiring the standard independence relation between the input and weak seed, i.e. that the mutual information $\mathrm{I}(X:Y)=0$, one can use randomness extractors in the presence of side-information that correlates the inputs by taking a penalty on the error or output length. The authors prove their results in the general cases of multi-source extractors, so we restate their theorems in the special case of two-source extractors.
In the case of classical side-information $E$, with $\mathrm{I}(X:Y|E)=0$ instead of $\mathrm{I}(X:Y)=0$ and with conditional min-entropy’s $\mathrm{H}_{\infty}(X|E)\geq k_{1}$ and $\mathrm{H}_{\infty}(Y|E)\geq k_{2}$;
@@
Any $(k_{1},k_{2},\epsilon,n_{1},n_{2},m)$ (strong) two-source extractor is a $(k_{1}-\log_{2}(\frac{1}{\epsilon}),k_{2}-\log_{2}(\frac{1}{\epsilon}),3\epsilon,n_{1},n_{2},m)$ classical-proof randomness extractor in the Markov model.
In the quantum case (where the quantum side-information of the adversary is now a quantum system $\mathcal{E}$) one requires $\mathrm{I}(X:Y|\mathcal{E})=0$ instead of $\mathrm{I}(X:Y)=0$ and that both sources have conditional min-entropy, i.e. $\mathrm{H}_{\infty}(X|\mathcal{E})\geq k_{1}$ and $\mathrm{H}_{\infty}(Y|\mathcal{E})\geq k_{2}$.
@@
Any $(k_{1},k_{2},\epsilon,n_{1},n_{2},m)$ (strong) two-source extractor is a $(k_{1}-\log_{2}(\frac{1}{\epsilon}),k_{2}-\log_{2}(\frac{1}{\epsilon}),\sqrt{3\epsilon 2^{m-2}},n_{1},n_{2},m)$ quantum-proof randomness extractor in the Markov model.
@@
The Markov model is particularly relevant in protocols for randomness amplification [37, 16] and in quantum key distribution (QKD) [^7] In quantum key distribution (QKD), if the adversary has some predictive power over the seed that is used in the privacy amplification step, one could also imagine that its actions on the quantum channel correlates the raw key material and the (now weak) seed. Such situations, in which the different sources becomes correlated through a third variable or system, can be accounted for by working in the Markov model.. Other works have considered different forms of quantum side-information, for example, where the adversary is allowed specific operations in order to build the side-information $\mathcal{E}$ [38] or when there is a bound on the dimension of the adversary’s side-information [39]. However, most of these other forms of side-information can be viewed as a particular case of the Markov one, making it a useful generic extension to use – see the discussion in [11] (Section 1.2: ‘Related work’). In the classical case, [10] explores what happens to the security of two-source extractors when the extractor input and the weak seed can be correlated in some specific models, for example, when they have bounded mutual information $I(X:Y)\leq t$, for some constant $t$.
@@ 3.2.3 Working with near-perfect seeds and smooth min-entropy bounds
Theorem 5.
[Extractors with (near-)perfect seeds, Appendix A in [40]] Given an $\epsilon_{s}$-perfect seed $Y$, which is used to extract $t$ times from a weak source [^8] Called a block min-entropy condition. The theorem can be generalised to the case that the min-entropy of each block differs, i.e. $\mathrm{H}_{\infty}(X_{i}|X_{i-1},...,X_{1})\geq k_{i}$. with $X_{i}$ for $i=1,...,t$ such that $\mathrm{H}_{\infty}(X_{i}|X_{i-1},...,X_{1})\geq k$ for all $i$ using a strong $(k,n_{2},\epsilon,n_{1},n_{2},m)$-seeded extractor $\mathsf{Ext_{s}}(X,Y)$ each time, the concatenated output $[\mathsf{Ext_{s}}(X_{1},Y),...,\mathsf{Ext_{s}}(X_{t},Y)]$ is $\epsilon_{t}$-perfect with
@@
Given a (strong) quantum-proof seeded extractor $(k,n_{2},\epsilon,n_{1},n_{2},m)$, one can use a weak input $X$ with smooth min-entropy $\mathrm{H}_{\infty}^{\delta}(X|\mathcal{E})\geq k$ at the cost of an additive error $\epsilon+\delta$, i.e. $(k,n_{2},\epsilon+\delta,n_{1},n_{2},m)$.
@@
Given a (strong) quantum-proof two-source extractor $(k_{1}-\log_{2}(1/\epsilon_{1})-1,k_{2}-\log_{2}(1/\epsilon_{2})-1,\epsilon,n_{1},n_{2},m)$ in the Markov model (see Section 3.2.2), one can use weak inputs $X$ and $Y$ satisfying $\mathrm{I}(X:Y|\mathcal{E})=0$ with smooth min-entropy’s $\mathrm{H}_{\infty}^{\delta_{1}}(X|\mathcal{E})\geq k^{\prime}_{1}$ and $\mathrm{H}_{\infty}^{\delta_{2}}(Y|\mathcal{E})\geq k^{\prime}_{2}$ giving a
$(k^{\prime}_{1}-\log_{2}(\frac{1}{\epsilon_{1}}),k^{\prime}_{2}-\log_{2}(\frac{1}{\epsilon_{2}}),2\epsilon+6(\delta_{1}+\delta_{2})+2(\epsilon_{1}+\epsilon_{2}),n_{1},n_{2},m)\mathrm{-extractor}$ (13)
This result also applies to classical-proof extractors and, because of the generality of the Markov model, to many other forms of (quantum) side-information [11].
3.2.4 Extractor extensions to more settings
As some of our extractors require specific input lengths (for example, that the input length must be prime), we give some results that give flexibility to increase or decrease the length of the extractor input, which later allow us to optimise the output length. First, we show that any two-source extractor is a seeded extractor, using only a short seed padded with fixed bits (i.e. these additional bits do not have any entropy).
Theorem 8 (Short seeded extractors from two-source ones).
A $(k_{1},k_{2},\epsilon,n_{1},n_{2},m)$-two-source randomness extractor $\mathsf{Ext_{2}}$ is also a $(k_{1},k_{2},\epsilon,n_{1},k_{2},m)$-seeded randomness extractor $\mathsf{Ext_{s}}$, where $\mathsf{Ext_{s}}(X,Y)=\mathsf{Ext_{2}}(X,[Y,c])$ and $[Y,c]$ is the seed $Y\in\{0,1\}^{\lfloor k_{2}\rfloor}$ concatenated with a fixed string $c\in\{0,1\}^{n_{2}-\lfloor k_{2}\rfloor}$, e.g. $c=\{0\}^{n_{2}-\lfloor k_{2}\rfloor}$.
@@ 3.2.4 Extractor extensions to more settings
Let $Y\in\{0,1\}^{\lfloor k_{2}\rfloor}$ be the seed of the extractor $\mathsf{Ext_{s}}$, i.e. $\mathrm{H}_{\infty}(Y)=\lfloor k_{2}\rfloor$, and $c\in\{0,1\}^{n_{2}-\lfloor k_{2}\rfloor}$ a constant (or fixed) string, i.e. (informally) $\mathrm{H}_{\infty}(c)=0$. Then $Y^{\prime}\equiv[Y,c]$ has $\mathrm{H}_{\infty}(Y^{\prime})=\mathrm{H}_{\infty}([Y,c])=\mathrm{H}_{\infty}(Y)=\lfloor k_{2}\rfloor$ and is therefore a valid input to the two-source extractor. ∎
This theorem is useful in some cases, for example, if a user has access to a seed that is insufficiently large to extract from a given weak input string, or, for certain lengths of the weak input and extractors, one can obtain more advantageous parameters by ‘enlarging’ the seed in this way, rather than shortening it (as shown is an option in Lemma 1). For example, the $\mathsf{Circulant}$ extractor requires that the weak input length plus one must be a prime, and the closest larger prime to an arbitrary integer might be closer than the closest smaller one. We now show how to shorten the weak input size, if desired.
@@
A random variable $X\in\{0,1\}^{n}$ with min-entropy $\mathrm{H}_{\infty}(X)=k$ can be shortened by some integer $c$, $11\leq c\leq n-$ 1, to $X^{\prime}=X_{0:n-c-1}\in\{0,1\}^{n-c}$, whilst keeping min-entropy at-least $\mathrm{H}_{\infty}(X^{\prime})\geq k-c$.
@@
We have that
$\mathrm{H}_{\infty}(X^{\prime})$ $=-\log_{2}(\max\limits_{x^{\prime}}p_{X^{\prime}}(x^{\prime}))=-\log_{2}\left(\max\limits_{x^{\prime}}\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}p_{X^{\prime},\bar{X}^{\prime}}(x^{\prime},\bar{x}^{\prime})\right)\geq-\log_{2}\left(\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}\max\limits_{x^{\prime}}p_{X^{\prime},\bar{X}^{\prime}}(x^{\prime},\bar{x}^{\prime})\right)$ (14)
$\geq-\log_{2}\left(\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}\max\limits_{x^{\prime},\bar{x}^{\prime}}p_{X^{\prime},\bar{X}^{\prime}}(x^{\prime},\bar{x}^{\prime})\right)\geq-\log_{2}\left(\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}2^{-k}\right)=-\log_{2}(2^{-k}2^{c})=k-c$ (15)
where we have used $\bar{X}^{\prime}=X_{n-c:n-1}$ so that $[X^{\prime},\bar{X}^{\prime}]=X$. ∎

### v3

@@ 3.2.1 Two-source extractors in the product source model
Any strong classical-proof (or quantum-proof) $(n_{1},k_{1},n_{2},k_{2},m,\epsilon)$-seeded extractor is a strong classical-proof (or quantum-proof) $(n_{1},k_{1},n_{2},k_{2},m,2^{n_{2}-k_{2}}\epsilon)$-two-source extractor, strong in the (now) weak seed. If the extractor is constructed from a two-universal family of hash functions, it is a strong classical-proof (or quantum-proof) $(n_{1},k_{1},n_{2},k_{2},m,2^{(n_{2}-k_{2})/2}\epsilon)$-two-source extractor.
The quantum-proof claim in Theorem 1 does not directly cover quantum side information about the (now) weak seed, only on the weak input, and still demands that the weak seed and weak input are independent. Using the proof techniques of [46] (‘Proof of proposition 1’) this can be directly generalised to the case that an adversary has quantum side information about both the input and the weak seed, but with the constraint that, for extractors based on two-universal hash families, this side information is in product form. More concretely, this is the case when the input and the now weak seed form a so-called classical-classical quantum-quantum state, whereby the combined two-universal hashing based extractor input (for the input and (weak) seed given by Equation 4) takes the form
@@
This model for quantum-proof extractors is the product source model.
Theorem 2 (Quantum-proof two-source extension in the product source model, Theorem 1 with Proposition 1 in [46]).
Any strong quantum-proof $(n_{1},k_{1},n_{2},k_{2}=n_{2},m,\epsilon)$-seeded extractor based on two-universal hash families is a strong quantum-proof $(n_{1},k_{1},n_{2},k_{2},m,2^{(n_{2}-k_{2})/2}\epsilon)$-two-source extractor in the product source model.
@@
The proof of this follows directly from [46] ‘Proof of proposition 1’ by noticing that n1n-1 can be generically replaced with $n_{2}$, the length of the (weak) seed and that their bound on ()(*) (as appearing in the text) is found by bounding the collision entropy of the seeded version of the extractor, which holds for all extractors based off two-universal families of hash functions. ∎
@@ 3.2.2 Two-source extractors in the Markov model
Two-source extractors can also be made secure in the so-called Markov model [13], which considers stronger adversaries than in the product source model. Instead of requiring the standard independence relation between the input and weak seed, i.e. that the mutual information $\mathrm{I}(X:Y)=0$, one can use randomness extractors in the presence of side information that correlates the inputs by taking a penalty on the error or output length. The authors prove their results in the general cases of multi-source extractors, so we restate their theorems in the special case of two-source extractors.
In the case of classical side information $E$, with $\mathrm{I}(X:Y|E)=0$ instead of $\mathrm{I}(X:Y)=0$ and with conditional min-entropy’s $\mathrm{H}_{\infty}(X|E)\geq k_{1}$ and $\mathrm{H}_{\infty}(Y|E)\geq k_{2}$;
@@
Any (strong) $(n_{1},k_{1},n_{2},k_{2},m,\epsilon)$-two-source extractor is a (strong) classical-proof $(n_{1},k_{1}-\log_{2}(1/\epsilon),n_{2},k_{2}-\log_{2}(1/\epsilon),m,3\epsilon)$-two-source randomness extractor in the Markov model.
In the quantum case (where the quantum side information of the adversary is now a quantum system $\mathcal{E}$) one requires $\mathrm{I}(X:Y|\mathcal{E})=0$ instead of $\mathrm{I}(X:Y)=0$ and that both sources have conditional min-entropy, i.e. $\mathrm{H}_{\infty}(X|\mathcal{E})\geq k_{1}$ and $\mathrm{H}_{\infty}(Y|\mathcal{E})\geq k_{2}$.
@@
Any (strong) $(n_{1},k_{1},n_{2},k_{2},m,\epsilon)$-two-source extractor is a (strong) quantum-proof $(n_{1},k_{1}-\log_{2}(1/\epsilon),n_{2},k_{2}-\log_{2}(1/\epsilon),m,\sqrt{3\epsilon 2^{m-2}})$-two-source randomness extractor in the Markov model.
We now show that quantum-proof seeded extractors that are based on two-universal hash families can be transformed into two-source extractors in the Markov model with minimal penalty (only due to having a second weak source). This applies to both our Circulant construction and Toeplitz.
Corollary 1.
Any (strong) quantum-proof $(n_{1},k_{1},n_{2},k_{2}=n_{2},m,\epsilon)$-seeded extractor based on two-universal hash families is also a (strong) quantum-proof $(n_{1},k_{1},n_{2},k_{2},m,2^{(n_{2}-k_{2})/2}\epsilon)$-two-source extractor in the Markov model.
Proof.
Via Theorem 2, the quantum-proof $(n_{1},k_{1},n_{2},k_{2}=n_{2},m,\epsilon)$-seeded extractor becomes a quantum-proof $(n_{1},k_{1},n_{2},k_{2},m,2^{(n_{2}-k_{2})/2}\epsilon)$ two-source extractor in the product source model, as it is based on two-universal hash families. According to [47], a Markov state $\rho_{XY\mathcal{E}}$ (for two sources $X$, $Y$, and the adversary’s system $\mathcal{E}$, with the Markov conditions discussed in Theorem 4) can be decomposed into product states (Equation 11) as
$\rho_{XY\mathcal{E}}=\bigoplus_{t}p(t)\rho^{t}_{X\mathcal{E}_{X}^{t}}\otimes\rho^{t}_{Y\mathcal{E}_{Y}^{t}}$ (12)
where $t$ runs over a finite alphabet with probability distribution $p(t)$, and $\mathcal{H}_{\mathcal{E}}=\bigoplus_{t}\mathcal{H}_{\mathcal{E}_{X}^{t}}\otimes\mathcal{H}_{\mathcal{E}_{Y}^{t}}$ is the Hilbert space of $\mathcal{E}$ (the quantum system held by the adversary). Using this decomposition, we can rewrite Equation 6 as
$\left\|\rho_{\text{Ext}(X,Y)\mathcal{E}}-u_{\mathcal{Z}}\otimes\rho_{\mathcal{E}}\right\|_{1}$ $=\left\|\text{Ext}\otimes\mathbb{I}_{\mathcal{E}}\left(\rho_{XY\mathcal{E}}\right)-u_{\mathcal{Z}}\otimes\rho_{\mathcal{E}}\right\|_{1}$ (13)
$=\left\|\text{Ext}\otimes\mathbb{I}_{\mathcal{E}}\left(\bigoplus_{t}p(t)\rho^{t}_{X\mathcal{E}_{X}^{t}}\otimes\rho^{t}_{Y\mathcal{E}_{Y}^{t}}\right)-u_{\mathcal{Z}}\otimes\left(\bigoplus_{t}\rho^{t}_{\mathcal{E}_{X}^{t}}\otimes\rho^{t}_{\mathcal{E}_{Y}^{t}}\right)\right\|_{1}$ (14)
$=\sum_{t}p(t)\left\|\text{Ext}\otimes\mathbb{I}_{\mathcal{E}}\left(\rho^{t}_{X\mathcal{E}_{X}^{t}}\otimes\rho^{t}_{Y\mathcal{E}_{Y}^{t}}\right)-u_{\mathcal{Z}}\otimes\rho^{t}_{\mathcal{E}_{X}^{t}}\otimes\rho^{t}_{\mathcal{E}_{Y}^{t}}\right\|_{1}$ (15)
$\leq\sum_{t}p(t)2^{(n_{2}-k_{2})/2}\epsilon=2^{(n_{2}-k_{2})/2}\epsilon$ (16)
where $\text{Ext}(X,Y)$ denotes the application of the extractor on $X,Y$, which is also understood as the CPTP map $\text{Ext}\otimes\mathbb{I}_{\mathcal{E}}$ on $\rho_{XY\mathcal{E}}$, and $u_{\mathcal{Z}}$ is the maximally mixed state over the output $Z$ of the extractor. In the last inequality, we have used the fact that the extractor is already quantum-proof in the product source model, meaning that, $\left\|\text{Ext}\otimes\mathbb{I}_{\mathcal{E}}\left(\rho^{t}_{X\mathcal{E}_{X}^{t}}\otimes\rho^{t}_{Y\mathcal{E}_{Y}^{t}}\right)-u_{\mathcal{Z}}\otimes\rho^{t}_{\mathcal{E}_{X}^{t}}\otimes\rho^{t}_{\mathcal{E}_{Y}^{t}}\right\|_{1}\leq 2^{(n_{2}-k_{2})/2}\epsilon$ for all $t$, noting that the ccq-state $\rho^{t}_{X\mathcal{E}_{X}^{t}}\otimes\rho^{t}_{Y\mathcal{E}_{Y}^{t}}$ satisfies $\mathrm{H}_{\infty}(X|\mathcal{E})\geq k_{1}$ and $\mathrm{H}_{\infty}(Y|\mathcal{E})$ for any $t$, given $\mathrm{H}_{\infty}(X|\mathcal{E})\geq k_{1}$ and $\mathrm{H}_{\infty}(Y|\mathcal{E})$ for $\rho_{XY\mathcal{E}}$. ∎
In both Theorems 3, 4 and Corollary 1, one has the option to lift an extractor to the Markov model whilst maintaining whether it is strong or not (a strong extractor will be strong in the Markov model, a weak one will remain weak). We have indicated this option by putting strong in parenthesis in the statements.
@@
The Markov model is particularly relevant in protocols for randomness amplification [48, 18] and in quantum key distribution (QKD) [^8] In quantum key distribution (QKD), if the adversary has some predictive power over the seed that is used in the privacy amplification step, one could also imagine that its actions on the quantum channel correlates the raw key material and the (now weak) seed. Such situations, in which the different sources become correlated through a third variable or system, can be accounted for by working in the Markov model.. Other works have considered different forms of quantum side information, for example, where the adversary is allowed specific operations to build the side information $\mathcal{E}$ [49] or when there is a bound on the dimension of the adversary’s side information [50]. However, most of these other forms of side information can be viewed as a particular case of the Markov one, making it a useful generic extension to use – see the discussion in [13] (Section 1.2: ‘Related work’). In the classical case, [12] explores what happens to the security of two-source extractors when the extractor input and the weak seed can be correlated in some specific models, for example, when they have bounded mutual information $I(X:Y)\leq t$, for some constant $t$.
@@ 3.2.3 Working with near-perfect seeds and smooth min-entropy bounds
Theorem 5 (Extractors with near-perfect seeds, Appendix A in [51]).
Given an $\epsilon_{s}$-perfect seed $Y$, which is used to extract $t$ times from a weak input [^9] Called a block min-entropy condition. The theorem can be generalised to the case that the min-entropy of each block differs, i.e. $\mathrm{H}_{\infty}(X_{i}|X_{i-1},...,X_{1})\geq k_{i}$. with $X_{i}$ for i=1,,ti=1,...,t such that $\mathrm{H}_{\infty}(X_{i}|X_{i-1},...,X_{1})\geq k$ for all $i$ using a strong $(n_{1},k_{1}=k,n_{2},k_{2}=n_{2},m,\epsilon)$-seeded extractor $\mathsf{Ext_{s}}(X,Y)$ each time, the concatenated output $\mathsf{Ext_{s}}(X_{1},Y),...,\mathsf{Ext_{s}}(X_{t},Y)]$ is $\epsilon_{t}$-perfect with
@@
Given a (strong) quantum-proof $(n_{1},k_{1}=k,n_{2},k_{2}=n_{2},m,\epsilon)$-seeded extractor, one can use a weak input $X$ with smooth min-entropy $\mathrm{H}_{\infty}^{\delta}(X|\mathcal{E})\geq k$ at the cost of an additive error $\epsilon+\delta$, i.e. giving a $(n_{1},k_{1}=k,n_{2},k_{2}=n_{2},m,\epsilon+\delta)$-seeded extractor.
@@
Given a (strong) quantum-proof $(n_{1},k_{1}-\log_{2}(1/\epsilon_{1})-1,n_{2},k_{2}-\log_{2}(1/\epsilon_{2})-1,m,\epsilon)$-two-source extractor in the Markov model (see Section 3.2.2), one can use weak inputs $X$ and $Y$ satisfying $\mathrm{I}(X:Y|\mathcal{E})=0$ with smooth min-entropy’s $\mathrm{H}_{\infty}^{\delta_{1}}(X|\mathcal{E})\geq k^{\prime}_{1}$ and $\mathrm{H}_{\infty}^{\delta_{2}}(Y|\mathcal{E})\geq k^{\prime}_{2}$ giving a
$\left(n_{1},k^{\prime}_{1}-\log_{2}\left(\frac{1}{\epsilon_{1}}\right),n_{2},k^{\prime}_{2}-\log_{2}\left(\frac{1}{\epsilon_{2}}\right),m,2\epsilon+6(\delta_{1}+\delta_{2})+2(\epsilon_{1}+\epsilon_{2})\right)\mathrm{-extractor}$ (18)
This result also applies to classical-proof extractors and, because of the generality of the Markov model, to many other forms of (quantum) side information [13].
3.2.4 Extending extractors to more settings
As some of our extractors require specific input lengths (for example, that the input length must be prime with primitive root 2), we give some results that give flexibility to increase or decrease the length of the extractor input, which later allows us to optimise the output length. First, we show that any two-source extractor can be used as a seeded extractor with a short seed, by concatenating the seed with fixed bits.
Theorem 8 (Short seeded extractors from two-source extractors).
Any $(n_{1},k_{1},n_{2},k_{2},m,\epsilon)$-two-source randomness extractor $\mathsf{Ext_{2}}$ is a $(n_{1},k_{1},n_{2}^{\prime}=k_{2},k_{2},m,\epsilon)$-seeded randomness extractor $\mathsf{Ext_{s}}$ for any $k_{2}\leq n_{2}$, where $\mathsf{Ext_{s}}$ is constructed by first concatenating the seed (of length $k_{2}$) with a fixed bit string of length $n_{2}-k_{2}$ and then applying $\mathsf{Ext_{2}}$. Specifically, given a seed $Y\in\{0,1\}^{k_{2}}$, the seeded extractor is constructed from the two-source one as $\mathsf{Ext_{s}}(X,Y)=\mathsf{Ext_{2}}(X,[Y,c])$, where $[Y,c]$ denotes the concatenation of the seed $Y$ with any fixed string $c\in\{0,1\}^{n_{2}-k_{2}}$ (e.g. $c=\{0\}^{n_{2}-k_{2}}$).
@@ 3.2.4 Extending extractors to more settings
Let $Y\in\{0,1\}^{k_{2}}$ be a seed (i.e. it has full min-entropy $\mathrm{H}_{\infty}(Y)=k_{2}$) and $c\in\{0,1\}^{n_{2}-k_{2}}$ be a fixed string, e.g. $\{0\}^{n_{2}-k_{2}}$, with $\mathrm{H}_{\infty}(c)=0$. Define $Y^{\prime}$ as the concatenation of $Y$ and $c$, $Y^{\prime}=[Y,c]$, which has min-entropy $\mathrm{H}_{\infty}(Y^{\prime})=\mathrm{H}_{\infty}([Y,c])=\mathrm{H}_{\infty}(Y)=k_{2}$. Since $Y^{\prime}$ is a random variable with min-entropy $k_{2}$ and length $n_{2}$, it can be used as the second input to any $(n_{1},k_{1},n_{2},k_{2},m,\epsilon)$-two-source extractor $\mathsf{Ext_{2}}(X,Y^{\prime})$. Therefore, $\mathsf{Ext_{2}}(X,Y^{\prime})$ can be understood as a seeded extractor $\mathsf{Ext_{s}}(X,Y)$, constructed by first concatenating the seed $Y$ with a fixed string of bits then applying $\mathsf{Ext_{2}}$. In other words, $\mathsf{Ext_{s}}(X,Y)=\mathsf{Ext_{2}}(X,[Y,c])=\mathsf{Ext_{2}}(X,Y^{\prime})$. ∎
This theorem is useful when a seed is too short to be used in a seeded extractor for a given weak input. For example, consider a seed $Y$ of length $n_{2}$ and a weak input $X\in\{0,1\}^{n_{1}}$ with $n_{1}>n_{2}$. If $n_{1}$ is sufficiently large compared to $n_{2}$, the seed length $n_{2}$ may be insufficient to perform extraction using existing seeded extractor constructions. However, by concatenating $Y$ with a zero string of length $c$, one can create a weak seed $Y^{\prime}$ of length $n_{2}+c$ with the same min-entropy $k_{2}=n_{2}$. This weak seed is now a valid second input for a two-source extractor, which may give more freedom and allow extraction to become possible.
We now show how to shorten the weak input length, if desired.
@@
For any integer $c\leq n$, a random variable $X\in\{0,1\}^{n}$ with min-entropy $\mathrm{H}_{\infty}(X)=k$ can be shortened to a new random variable $X^{\prime}\in\{0,1\}^{n-c}$ with min-entropy $\mathrm{H}_{\infty}(X^{\prime})\geq k-c$, by removing the last $c$ bits of $X$.
@@
Let $X^{\prime}\in\{0,1\}^{n_{1}-c}$, $\bar{X}^{\prime}\in\{0,1\}^{c}$ and $X=[X^{\prime},\bar{X}^{\prime}]\in\{0,1\}^{n_{1}}$, where $[,]$ denotes concatenation. By definition, $\mathrm{H}_{\infty}(X)=k$ and we want to prove a lower bound on $\mathrm{H}_{\infty}(X^{\prime})$. Using that $p_{X^{\prime}}(x^{\prime})$ is a marginal distribution of $p_{[X^{\prime},\bar{X}^{\prime}]}([x^{\prime},\bar{x}^{\prime}])$, we have that
$\mathrm{H}_{\infty}(X^{\prime})$ $=-\log_{2}(\max\limits_{x^{\prime}}p_{X^{\prime}}(x^{\prime}))=-\log_{2}\left(\max\limits_{x^{\prime}}\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}p_{[X^{\prime},\bar{X}^{\prime}]}([x^{\prime},\bar{x}^{\prime}])\right).$ (19)
Continuing from Equation 19 by passing the maximum inside the sum and then maximise over both variables (instead of one), we get
$-\log_{2}\left(\max\limits_{x^{\prime}}\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}p_{[X^{\prime},\bar{X}^{\prime}]}([x^{\prime},\bar{x}^{\prime}])\right)\geq-\log_{2}\left(\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}\max\limits_{x^{\prime}}p_{[X^{\prime},\bar{X}^{\prime}]}([x^{\prime},\bar{x}^{\prime}])\right)$ (20)
$\geq-\log_{2}\left(\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}\max\limits_{[x^{\prime},\bar{x}^{\prime}]}p_{[X^{\prime},\bar{X}^{\prime}]}([x^{\prime},\bar{x}^{\prime}])\right)=-\log_{2}\left(\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}\max\limits_{x}p_{X}(x)\right)$ (21)
$\geq-\log_{2}\left(\sum\limits_{\bar{x}^{\prime}\in\{0,1\}^{c}}2^{-k}\right)=-\log_{2}(2^{-k}2^{c})=k-c$ (22)
where we have used that $\max\limits_{x}p_{X}(x)\leq 2^{-k}$, since $\mathrm{H}_{\infty}(X)=k$ by definition. ∎